Find the slope-intercept form of the line passing through the point (-1, -8) and parallel to the line 4x + 3y =2.
Answer by nerdybill(7384)

Find the slope-intercept form of the line passing through the point (-1, -8) and parallel to the line 4x + 3y =2.
.
determine slope of:
4x + 3y =2
by putting it into the "slope-intercept" form:
4x + 3y =2
3y = -4x+2
y = (-4/3)x + 2/3
slope then is -4/3
.
A line that is parallel must have the same slope (-4/3). Use that and the given point (-1, -8) to plug into the "point-slope" form:
y - y1 = m(x - x1)
y - (-8) = (-4/3)(x - (-1))
y + 8 = (-4/3)(x + 1)
y + 8 = (-4/3)x - 4/3
y = (-4/3)x - 4/3 - 8
y = (-4/3)x - 4/3 - 24/3
y = (-4/3)x - 28/3 (this is what they're looking for)